OpenStudy (anonymous):

Hi, Is it possible (and if it is, how) to find the equation of a function with absolute values when I know a graph of that function? I mean functions like f(x) = |x-3| + |2-5x|

OpenStudy (lgbasallote):

....so you're looking to find the equation of a function......given an equation of a function....?

OpenStudy (anonymous):

I only know a graph.

OpenStudy (lgbasallote):

..but you gave an equation....

OpenStudy (amistre64):

f(x) = g(x) + h(x) 

OpenStudy (lgbasallote):

^he/she probably made more sense than i did

OpenStudy (amistre64):

lol, im a he/she ??

OpenStudy (lgbasallote):

sorry...i forgot....."it" i just always get the feeling you're very close to human-ness so i call you a he/she

OpenStudy (amistre64):

us similated answering programs have rights too ;)

OpenStudy (amistre64):

this would be easier if we had a graph to go by

OpenStudy (anonymous):

In particular how to find the equation of

OpenStudy (anonymous):

I know it can be written as a picewise function , but it is possible to write it as equation with absolute values.

OpenStudy (amistre64):

the vertexes might be able to give us a clue as to that

OpenStudy (amistre64):

-1 = g(1) + h(1) 1 = g(0) + h(0) 5 = g(2.5) + h(2.5) 5 = g(3) + h(3)

OpenStudy (anonymous):

OK but the equation might contain more than two absolute values in it.

OpenStudy (amistre64):

yeah, i was thinking of that as well; but this is just a general idea

OpenStudy (amistre64):

im wondering if we can assume that the given slopes are the averages parts

OpenStudy (amistre64):

also the absolute values need not be linear either

OpenStudy (amistre64):

but it prolly helps if they are linear http://www.wolframalpha.com/input/?i=y%3D+%7Cx%5E2-6%7C

OpenStudy (anonymous):

Equation might be in the form: \[ f(x) = a_1x + b_1 + a_2|x + b_2| + a_3|x + b_3| + \ldots \]

OpenStudy (anonymous):

where \[a_1,a_2, a_3, \ldots, b_1, b_2, b_3, \ldots \] can be any real numbers

OpenStudy (amistre64):

im curious if we can conform this by using a method of zeroing out coeffs

OpenStudy (amistre64):

\[y = c_1~x|x-1||x-2.5|+c_2~x|x-1||x-3|+c_3~x|x-2.5||x-3|\]\[\hspace{3em}+c_4~|x-1||x-2.5||x-3|\]

OpenStudy (amistre64):

\[f(x)\quad \begin{matrix}-2x+1~;~x\in(-\infty,1)\\4x-5~;~x\in \left[1,\frac52 \right)\\5~;~x\in \left[\frac52,3 \right] \\2x-1~;~x\in (3,\infty)\end{matrix}\] i wonder, since there are 4 piecewise parts, if this could be comprised of 4 absolutes value function

OpenStudy (anonymous):

I think, that it can be written as a sum of 3 functions of a form a|x + b|

OpenStudy (amistre64):

do you have material that goes with this?

OpenStudy (amistre64):

if its just some idea that you are working with, thats fine too; but if there is material associated with it, that could help to focus the thoughts a little :)

OpenStudy (anonymous):

No, I don't. I'm just learning about absolute value functions and i do some exercises in a book for graphing. I was curious if i can reverse the "process".

OpenStudy (amistre64):

im thinking that since we have 4 different slopes on the graph, that we would need to average 4 different points in each interval to match. just an idea. what is your notion for it being 3?

OpenStudy (anonymous):

Just messing around with GeoGebra. When it has 3 absolute values it has 3 breaks.

OpenStudy (amistre64):

3 breaks yes, but i cant see a way for one of those breaks to be horizontal

OpenStudy (amistre64):

y = |x+1| + |x-1| gives us a horizontal; anytime we provide a shift, we can get a horizontal section

OpenStudy (anonymous):

It looks that there is a horizontal part in some interval if sum of two absolute value functions have opposite slope in that interval.

OpenStudy (anonymous):

Or sum of their slopes is zero.

OpenStudy (anonymous):

In general slope of the function in some interval is equal to sum of slopes of functions in that interval.

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y%3D2%282%7Cx-1%7C-1+%2B+4%7Cx-5%2F2%7C%2B5+-+2%7Cx-3%7C%2B5%29-19 that looks promising if you can work out the dynamics between them

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y%3D3%7Cx-1%7C-1+-2%7Cx-5%2F2%7C%2B3%2B%7Cx-3%7C-2 i cant tell, but this is sooo close

OpenStudy (amistre64):

i wish i knew what i did, but i kept working with the vertices somehow

OpenStudy (anonymous):

I made some progress. Assume that the function consists of sum of 3 functions of a form a|x + b| So it will be \[ f(x) = a|x + b| + c|x + d| + d|x+e| \] where a,b,c,d,e are real numbers. Function has breaks at points 1, 2.5 and 3, hence \[ f(x) = a|x - 1| + c|x - 2.5| + d|x-3| \] And then I can divide the function into 4 intervals Let them denote \[ I_1 = (-\infty, 1) \\ I_2 = [1, 2.5) \\ I_3 = [2.5, 3) \\ I_4 = [3, +\infty) \] As I discovered before, slope of the function is the sum of slopes of function the function is build from. So I can solve for slope in each interval. Thaen I get system of equations \[ a + c + e = 2 \\ a- c - e = 4 \\ a + c - e = 0 \] From which I get \[ a = 3 \\ c = -2 \\ e = 1\] And therefore the function is \[\Large f(x) = 3|x-1| - 2|x-2.5| + |x-3| \] http://www.wolframalpha.com/input/?i=+f%28x%29+%3D+3 |x-1|+-+2|x-2.5|++%2B+|x-3|

OpenStudy (amistre64):

that seems more systematic than the approach i took -- some skillm and alot of luck ;)

OpenStudy (amistre64):

how did you arrive at this setup ?? a+c+e=2 a−c−e=4 a+c−e=0

OpenStudy (amistre64):

also you have a 4th slope, how did you figure which one to ignore?

OpenStudy (anonymous):

the fourth one was -(a+c+e) = -2 which is -1 multiple of the first

OpenStudy (anonymous):

Ok in the first interval the equation is \[ -a|x - 1| - c|x- 2.5| - e|x - 3| = -(a + c + e)x + a + 2.5c + 3e\] but i only need a slope which is \[ -(a+c+e) \] from the graph you can see that slope in that interval is -2 hence the first equation is \[ -(a+c+e) = -2 \] Slopes in other intervals are done in the same way

OpenStudy (anonymous):

Oh, in the previous post the equation should be without absolute values \[ −a(x−1)−c(x−2.5)−e(x−3)=−(a+c+e)x+a+2.5c+3e \]

OpenStudy (amistre64):

a|x-1| +b|x-2.5| +c|x-3| = -2 a|x-1| +b|x-2.5| +c|x-3| = 4 a|x-1| +b|x-2.5| +c|x-3| = 0 a|x-1| +b|x-2.5| +c|x-3| = 2 i still cant see how you get from that to the setup a+c+e=2 a−c−e=4 a+c−e=0

OpenStudy (amistre64):

im glad you did of course :) but i cant see a reasoning to it

OpenStudy (amistre64):

im using abc for your ace of course

OpenStudy (anonymous):

In the first interval expression in all three absolute values is negative so you can write the equation as \[ −a(x−1)−c(x−2.5)−e(x−3) \] slope is -(a+c+e) in the second interval is x- 1 non negative and x - 2.5 and x- 3 are negative so it can be written as \[a(x-1) - c(x-2.5) - e(x-3) \] slope is a-c-e In the third interval x-1 and x-2.5 are non-negative and x-3 is negative therefore it can be written as \[ a(x−1)+c(x−2.5)−e(x−3) \] slope is a+c-e Last, in the fourth equation are all expression in absolute value non-negative, and can be written as \[ a(x−1)+c(x−2.5)+e(x−3) \] slope is a+c+e From which i get the system of equations \[ -(a+c+e) = -2 \\ a-c-e = 4 \\ a+c-e = 0 \\ a+c+e = 2 \]

OpenStudy (amistre64):

i would conjecture from that, that we start out assuming that the slopes are all equal, then switch them out one by one till, in the different intervals we have the slope of one of them flipping ....

OpenStudy (amistre64):

equal meaning of course in direction + or -

OpenStudy (amistre64):

spose we had 5 slopes, such that the ends are "equal" but opposite then: a+b+c+d = s1 a+b+c-d = s2 a+b-c-d = s3 a-b-c-d = s4 is that a fair idea?

OpenStudy (amistre64):

i wonder if we can construct a summation of these things such that the ends have different slopes ....

OpenStudy (anonymous):

The method of division into intervals when you solve an equation and solve in each interval is something I learn in school and I've never seen in English book or website. Is that method taught in English-speaking countries? If not I can try to explain it in more detail.

OpenStudy (amistre64):

im not sure what is in the schools these days; ive been out of class for 25 years or so :)

OpenStudy (amistre64):

if the ends have different slopes; say -2 on the left and 3 on the right, then you cant construct it with absolutes. If both ends have the same slope that differ by a sign; then it appears you can

OpenStudy (amistre64):

if you use an extra third vertex at say (inf,inf) maybe :)

OpenStudy (anonymous):

Then it probably has \[ a_1 \neq 0, \\ b_1 \neq 0 \] in the equation [\ f(x)=a_1x+b_1+a_2|x+b_2|+a_3|x+b_3|+\ldots \]

OpenStudy (amistre64):

your first delimiter is flipped about, spose to be \[

OpenStudy (anonymous):

\[ f(x)=a_1x+b_1+a_2|x+b_2|+a_3|x+b_3|+\ldots \]

OpenStudy (amistre64):

ah yes, that plain old linear in the front does help to get different slopes

OpenStudy (amistre64):

as long as a1 not= 0, i can form different slopes; the b1 is superfluous since that only adjusts the height overall and not the slope